OpenStudy (anonymous):

"Find the vertex of the graph of the function" f(x)=(x+6)^2-5

OpenStudy (anonymous):

Can you simplify it into f(x)=Ax+By+C form?

OpenStudy (anonymous):

*Ax^2+Bx+C, I was thinking of standard form for some reason

OpenStudy (anonymous):

how would that be finding the vertex

OpenStudy (anonymous):

?

jigglypuff314 (jigglypuff314):

the equation is already in the vertex form for parabolas therefore the vertex of your equation of f(x)=(x+6)^2-5 is (-6, 5) because when f(x)=a(x-h)^2-k, (h, k) is the vertex hope this helped :)